OpenStudy (anonymous):

if f(x) =x+5 and g(x) -1/(x+1) find f(g(1/2))

OpenStudy (anonymous):

with explanation please

OpenStudy (ghazi):

put x=(-1/x+1) in f(x) to get f(g(x))

OpenStudy (ghazi):

your f(g(x))= (1/x+1)+5\[\frac{ -1 }{ x+5 }+5=\frac{ -1+5(x+5) }{ x+5 }=\frac{ -1+5x+25 }{ x+5 }=\frac{ 5x-24 }{ x+5 }\]

OpenStudy (ghazi):

now put x=1/2 to get answer